HUBBLE SPACE TELESCOPE DAILY REPORT #1368 PERIOD COVERED: 0000Z (UTC) 04/12/95 - 0000Z (UTC) 04/13/95 Daily Status Report as of 103/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 5397 (Determination of the Extragalactic Distance Scale: Cycle 4 High) The WF/PC-2 was used to make images of galaxies NGC-4725 and NGC-4414 to be used in measuring the periods of Cepheid variable stars within the galaxies. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Three Sets of FOS/RD 5535 (Cycle 4 Internal Wavelength Calibration) The Faint Object Spectrograph (Red Side) was used to make internal calibration observations to monitor the stability of the wavelength scale for different gratings. The observations were completed as planned, and no problems were reported. 1.3 Completed Two Sets of GHRS/2 5601 (GHRS ECH-A Wavelength Monitor: Cycle 4) The High Resolution Spectrograph was used to make Echelle-B spectral calibration observations using lamp #2. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6143 (WF/PC-2 Cycle 4 Photometric Calibration: Smaller Set, Part II) The WF/PC-2 was used to make single CCD calibration observations of standard star GRW+70D582 to monitor the photometric stability and quantum efficiency of the PC1 and WF3. The observations were executed as scheduled, and no anomalies were reported. 1.5 Completed WF/PC-2 5629 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1C: 4-Chip STD in F170W) The WF/PC-2 was used to make photometric calibration observations of the ultraviolet standard star GRW+70D582. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 5672 (Ages For The Outermost Globular Clusters) The WF/PC-2 was used to make both deep and short exposure observations of the central regions of globular cluster PAL3-Centaurus. This cluster was chosen as it represents an early staging in the formation of the galaxy. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Significant HSTARs: None 2.4 Operations Notes: The STGT data inversions continue, as do the bit slips while dumping ESTR #2. Due to problems with the UPS, the backup power system, both the ESS and TALOS were unavailable for slightly over eight hours last night. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-